OpenStudy (7thgrader):

Solve the equation. Describe the solution to the equation as one solution, no solution, or an infinite number of solutions. If there is one solution, then include the value for the variable. Explain your reasoning. Show all your work.

OpenStudy (anonymous):

\(\Huge\color{Orange}{\rightarrow}\color{Orange}{\rightarrow}\cal\color{Blue}{\bigstar I'll~Try\bigstar}\color{Orange}{\leftarrow}\color{Orange}{\leftarrow}\)

OpenStudy (7thgrader):

1/8 (x + 24) = 5/8x + 9 -1/2x

OpenStudy (7thgrader):

@JohnKnows

OpenStudy (7thgrader):

@Gatorgirl @math&ing001 @hotdancingbabe14 @LOLAri @TheRaggedyDoctor @texaschic101

OpenStudy (math&ing001):

8 * [1/8 (x + 24) = 5/8x + 9 -1/2x] x + 24 = 5x + 72 - 4x x + 24 = x + 72 24 = 72

OpenStudy (7thgrader):

thanks :)

OpenStudy (math&ing001):

Welcome !